# Boaz Moerman

### v2

@@ 3.3. Examples of $\mathcal{M}$-points
In this section we give some examples of $\mathcal{M}$-points, such as integral points, Campana points and Darmon points, as well as their geometric counterparts. For more examples, see [Moe24, §3.3]. We fix a proper variety $X$ over a number field $K$ with an integral model $\mathcal{X}$ over $\mathcal{O}_{S}$ for some finite set of places $S\subset\Omega_{K}$, and we fix an $\mathcal{O}_{S}$-scheme $Y$ (such as the spectrum of $\mathcal{O}_{S}$ or $\mathcal{O}_{v}$ for a place $v\in\Omega_{K}\setminus S$). On $X$, we choose prime divisors $D_{1},\dots,D_{n}$ and we denote their Zariski closures in $\mathcal{X}$ by $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$.
@@
If $(\mathcal{X},\mathcal{M})$ is the pair given by $\mathfrak{M}=\{(0,\dots,0)\}$, then the $\mathcal{M}$-points over $R$ on the pair are exactly the integral points over $R$ on $\mathcal{U}=\mathcal{X}\setminus\bigcup_{i=1}^{n}\mathcal{D}_{i}$. More generally, if $(\mathcal{X},\mathcal{M})$ is the pair given by $\mathfrak{M}=\{\mathbf{m}\mid m_{i}=0\,\forall i\in I\}$ for $I\subset\{1,\dots n\}$, then the $\mathcal{M}$-points on the pair are the integral points on $\mathcal{X}\setminus\bigcup_{i\in I}\mathcal{D}_{i}$.
@@
Note that for $I\subset\{1,\dots,n\}$, the Darmon points an the (weak) Campana points on $(\mathcal{X},\sum_{i\in I}\mathcal{D}_{i})$ are just the integral points on $\mathcal{X}\setminus\bigcup_{i\in I}\mathcal{D}_{i}$ as in Example 3.22.
@@
As explained in [Moe24, §8], Darmon points can be viewed as integral points on the root stack $(\mathcal{X},\sqrt[\mathbf{m}]{\mathcal{D}})$ corresponding to the Campana pair $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$.
@@
Geometric Campana points can be viewed as the subset of Campana points which remain Campana points after enlarging the base field. In particular, geometric Campana points are just the same as Campana points whenever the support of the divisor $D_{\mathbf{m}}$ is a strict normal crossings divisor (over the ground field $K$). For Darmon points analogous statements hold. Let us consider an example where geometric Campana points differ from Campana points.
@@
Besides Campana points, Darmon points, and (weak) Campana points, there are many more interesting sets of $\mathcal{M}$-points, see [Moe24, §3.3].
@@ 4. The Picard group of a pair and its geometry
For pairs $(X,M)$ corresponding to Darmon points, then there is a natural definition for the Picard group of $(X,M)$. For such a pair, we have a corresponding root stack $(X,\sqrt[\mathbf{m}]{D})$ which lies over $X$. In this case we want the Picard group of $(X,M)$ to be the Picard group of $(X,\sqrt[\mathbf{m}]{D})$, and the canonical divisor class to be the canonical divisor class on the stack. Both the Picard group and the canonical divisor class can be computed using the natural morphism $(X,\sqrt[\mathbf{m}]{D})\rightarrow X$, which is an isomorphism over the open set $U=X\setminus(D_{1}\cup\dots\cup D_{n})$ and ramified over the divisors $D_{i}$ with multiplicity $m_{i}$.
More generally, we can regard $(X,M)$ as some sort of geometric space akin to a scheme or stack lying above $X$. If $(X,M)=(X,M_{\operatorname{mon}})$ then $(X,M)$ determines a functor $Y\mapsto(X,M)(Y)$ as we have seen in Remark 3.18. We view the natural inclusion map $(X,M)\rightarrow X$ as analogous to an “generically finite morphism” which is an isomorphism over $U$. Inspired by this view we introduce corresponding objects to the pair $(X,M)$, such as divisors and a Picard group.
@@ 4.1. Divisors on pairs
For a pair $(X,M)$ we define its set of generators to be the set $\Gamma_{M,\mathcal{C}}$ of all $(\mathbf{m},c)\in\mathfrak{M}_{\mathcal{C},\textnormal{fin}}$ such that the only subset $J\subset\mathfrak{M}_{\mathcal{C}}\setminus\{(\mathbf{0},X)\}$ such that
@@
is $J=\{(\mathbf{m},c)\}$. The subscript fin here indicates that we are only considering $(\mathbf{m},c)$ such that the coefficients of $\mathbf{m}$ are finite. If $(X,M)$ is a divisorial pair defined by a set of multiplicities $\mathfrak{M}\subset\overline{\mathbb{N}}^{n}$, then we set
@@
    every divisor $D_{i}$ is connected, nonempty and smooth over $K$ while $\sum_{i=1}^{n}D_{i}$ is a strict normal crossings divisor as defined in [Stacks, Tag 0BI9],
The use of the term “smooth” is motivated by the fact that strict normal crossings pairs as considered in logarithmic geometry are log smooth, see e.g. [Ogu18, Chapter IV, Example 3.1.14]. Furthermore, [Moe24, Proposition 8.3] shows that a root stack $(X,\sqrt[\mathbf{m}]{D})$ is smooth if its corresponding pair $(X,M)$ is smooth.
@@
where $c^{\prime}\subset c$. Then the pair $(X,M)$ defined using this data is equivalent to $(X,M)$. Furthermore, by the assumptions on $(X,M)$ it follows that $(X,M^{\prime})$ is smooth. ∎
@@
Let $D$ and $D^{\prime}$ be effective divisors on $X$ and let $(\mathbf{m},c)\in\mathbb{N}^{n}_{\mathcal{C}}$. We define
$\mu((\mathbf{m},c),D-D^{\prime}):=\gcd(\mathbf{m})\cdot v_{(\mathbf{m}/\gcd(\mathbf{m}),c)}\left(\frac{f}{f^{\prime}}\right),$
where $f,f^{\prime}\in\mathcal{O}_{X,c}$ are local equations for $D,D^{\prime}$.
@@
The pullback of divisors to $(X,M)$ is intimately related to the theory of weighted (stacky) blow-ups as described in [QuRy22]. For $(\mathbf{m},c)\in\mathbb{N}_{\mathcal{C}}^{n}$, consider an open subset $X^{\prime}\subset X$ such that the divisors $D_{1},\dots,D_{n}$ are principal on $X^{\prime}$ and $\tilde{D}_{\mathbf{m}}=\tilde{D}_{\mathbf{m},c}$. Let $I_{\bullet}$ be the graded $\mathcal{O}_{X^{\prime}}$-subalgebra of $\mathcal{O}_{X^{\prime}}[t]$ generated by $f_{1}t^{m_{1}},\dots,f_{n}t^{m_{n}}$, where $f_{i}=0$ is a local equation for $D_{i}$. If $\pi\colon\mathrm{Bl}_{I_{\bullet}}X^{\prime}\rightarrow X^{\prime}$ is the corresponding weighted blow-up as defined in [QuRy22, Definition 3.2.1] and $D\in\operatorname{Div}(X^{\prime})$, then the coefficient of the exceptional divisor in $\pi^{*}D$ is equal to $\mu((\mathbf{m},c),D)$.
@@ 4.2. The Picard group of a pair
If $(X,M)$ is a smooth split pair corresponding to Darmon points with associated root stack $(X,\sqrt[\mathbf{m}]{D})$ as in [Moe24, §8], then [Cad07, Corollary 3.12] implies that the map $\operatorname{Pic}(X,M)\rightarrow\operatorname{Pic}(X,\sqrt[\mathbf{m}]{D})$, given by sending $\tilde{D}_{m_{i}\mathbf{e}}$ to the Cartier divisor $\frac{1}{m_{i}}D_{i}$ for all $i\in\{1,\dots,n\}$ with $m_{i}<\infty$, is an isomorphism. Furthermore, this isomorphism is compatible with the pullback homomorphisms of $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,M)$ and $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,\sqrt[\mathbf{m}]{D})$.
@@
Since $(X,M)$ is proper, for each $i=1,\dots,n$ there exists an element $(d_{i}\mathbf{e}_{i},D_{i})\in\Gamma_{M,\mathcal{C}}$ for some integer $d_{i}$. In particular, the coefficient of $\tilde{D}_{d_{i}\mathbf{e}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D_{i}$ is $d_{i}$, while for all $j\in\{1,\dots,n\}$ different from $i$, the coefficient of $\tilde{D}_{\mathbf{m}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D_{j}$ is zero, as $D_{i}$ is not contained in $D_{j}$. Similarly, for all $D\in\operatorname{Div}(U)$, the the coefficient of $\tilde{D}_{d_{i}\mathbf{e}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D$ is also zero. Since, furthermore, the restriction of $\operatorname{pr}_{M}^{*}$ to $\operatorname{Div}(U)$ is injective, $\operatorname{pr}^{*}_{M}$ is indeed injective. This directly implies that the homomorphism on the Picard groups is also injective.
@@
Suppose that $D\in\operatorname{Div}(X,M)$ is an element whose image in $\operatorname{Pic}(X_{\overline{K}},M)$ is trivial. This implies that $D$ is the pullback of $\operatorname{div}f$ for $f\in\overline{K}(X)^{\times}$. The rational function $f$ is defined over some Galois extension $E/K$, and $\tilde{f}=\prod_{\sigma\in\operatorname{Gal}(E/K)}\sigma(f)\in K(X)^{\times}$. Since $D$ is left invariant under the Galois action it follows that the pullback of $\operatorname{div}(\sigma(f))$ to $(X_{E},M)$ is $D$ itself. Therefore, it follows that $\deg(E/K)D\in\operatorname{Div}(X,M)$ is a principal divisor, so $D]\in\operatorname{Pic}(X,M)$ is torsion, as desired. ∎
@@ 4.3. The canonical divisor and the effective cone
Thus the canonical class is defined in such a way that it satisfies an analogue of the Hurwitz formula for morphisms of curves (see e.g. [Har77, Chapter IV, Proposition 2.3]). If $(X,M)$ is a smooth pair corresponding to Darmon points on a Campana pair $(X,D_{\mathbf{m}})$, then the canonical divisor of $(X,M)$ agrees with the canonical divisor of the root stack $(X,\sqrt[\mathbf{m}]{D})$, see for example [VoZB22, Proposition 5.5.6] for the case when $X$ is a curve.
@@
The effective cone of a proper pair is strongly convex, as the next proposition shows.
@@
In fact, for computing the Fujita invariant and the $b$-invariant, we only need to consider small generators $(\mathbf{m},c)\in\Gamma_{M,\mathcal{C}}$, which often considerably simplifies computations. For a stratum $c$ we set
@@
Note that here it suffices to consider the smallest strata, since if $c\subset c^{\prime}$, then $P_{c^{\prime}}\subset P_{c}$. Furthermore, for a divisorial pair $(X,M)$, the pairs $(X,M^{\prime})$ and $(X,M^{\prime\prime})$ are both divisorial, with sets of multiplicities $\mathfrak{M}^{\prime}=\Gamma_{M}\cap\partial P$ and $\mathfrak{M}^{\prime\prime}=\Gamma_{M}\cap V(P)$, where $P$ is the polyhedron given as the convex hull of the set
@@
The proof is essentially the same argument as in Lemma 4.39 by using Lemma 4.21. ∎
@@ 5. Rationally connected pairs
In modern times, Manin’s Conjecture is often formulated for rationally connected varieties, see for example [LeSeKuSh22, Conjecture 1.2]. The notion of rationally connected varieties (see e.g. [KMM92, Definition-Remark 2.2]) has a natural extension to smooth proper pairs.
@@
A smooth proper pair $(X,M)$ over a field $K$ is rationally connected if there exists a nonempty open subvariety $V\subset X$ such that for each algebraically closed field $L/K$ and every two points $p_{1},p_{2}\in V(L)$, there exists a rational curve $C\subset X_{L}$ containing both points such that $C$ is the image of a morphism $f\in(X,M_{\operatorname{mon}})(\mathbb{P}^{1}_{L})$, where $(X,M_{\operatorname{mon}})$ is the pair from Notation 3.17.
@@
By spreading out [Poo17, Theorem 3.2.1], we can find a set of places $S^{\prime}$ containing $S$ such that $f\colon\mathbb{P}^{1}_{K}\rightarrow X$ lifts to a morphism $\tilde{f}\colon\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}\rightarrow\mathcal{X}$. The pullbacks $\tilde{f}^{*}\mathcal{D}_{1},\dots,\tilde{f}^{*}\mathcal{D}_{n}$ are effective divisors on $\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}$. By further enlarging $S^{\prime}$ if necessary, we can ensure that these divisors have no components supported above a prime $\mathfrak{p}$ in $\mathcal{O}_{S^{\prime}}$, which ensures that $\tilde{f}\in(\mathcal{X},\mathcal{M}_{\operatorname{mon}})(\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}})$. Let $P\in\mathbb{P}_{K}^{1}(K)$. By the valuative criterion of properness, $P$ corresponds to an unique integral point $\tilde{P}\colon\operatorname{Spec}\mathcal{O}_{S^{\prime}}\rightarrow\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}$. Now it follows from the functoriality discussed in Remark 3.18 that $\tilde{f}\circ\tilde{P}\in(\mathcal{X},\mathcal{M}_{\operatorname{mon}})(\mathcal{O}_{S^{\prime}})$. ∎
@@
Let $(X,M)$ be a smooth proper pair over a field $K$ such that $X$ is a toric variety and $D_{1},\dots,D_{n}$ are torus invariant prime divisors. Then $(X,M)$ is rationally connected.
@@
Without loss of generality, we can assume that $\mathfrak{M}$ is a monoid and that $X$ is a split toric variety (so that $(X,M)$ is a split pair). Let $L$ be an algebraically closed field containing $K$ and let $L(t)$ be the field of rational functions over $L$ in the variable $t$. Write $(X_{L(t)},M_{L(t)})$ for the pair over $L(t)$ obtained by base changing $X$ and $D_{1},\dots,D_{n}$ to $L(t)$. This pair has the obvious integral model $(X_{\mathbb{P}^{1}_{L}},M_{\mathbb{P}^{1}_{L}})$ over $\mathbb{P}^{1}_{L}$ given by base changing $(X,M)$ to $\mathbb{P}^{1}_{L}$. Now [Moe24, Theorem 1.3] implies that the pair $(X_{L(t)},M_{L(t)})$ satisfies $M_{L(t)}$-approximation off the place $(1:0)\in\mathbb{P}_{L}^{1}$. Consequently, the embedding

### v3

@@ 3.3. Examples of $\mathcal{M}$-points
In this section we give some examples of $\mathcal{M}$-points, such as integral points, Campana points and Darmon points, as well as their geometric counterparts. For more examples, see [30, §3.3]. We fix a proper variety $X$ over a number field $K$ with an integral model $\mathcal{X}$ over $\mathcal{O}_{S}$ for some finite set of places $S\subset\Omega_{K}$, and we fix an $\mathcal{O}_{S}$-scheme $Y$ (such as the spectrum of $\mathcal{O}_{S}$ or $\mathcal{O}_{v}$ for a place $v\in\Omega_{K}\setminus S$). On $X$, we choose prime divisors $D_{1},\dots,D_{n}$ and we denote their Zariski closures in $\mathcal{X}$ by $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$.
@@
If $(\mathcal{X},\mathcal{M})$ is the pair given by $\mathfrak{M}=\{(0,\dots,0)\}$, then the $\mathcal{M}$-points over $Y$ on the pair are exactly the integral points over $Y$ on $\mathcal{U}=\mathcal{X}\setminus\bigcup_{i=1}^{n}\mathcal{D}_{i}$. More generally, if $(\mathcal{X},\mathcal{M})$ is the pair given by $\mathfrak{M}=\{\mathbf{m}\mid m_{i}=0\,\forall i\in I\}$ for $I\subset\{1,\dots n\}$, then the $\mathcal{M}$-points on the pair are the integral points on $\mathcal{X}\setminus\bigcup_{i\in I}\mathcal{D}_{i}$.
@@
Note that for $I\subset\{1,\dots,n\}$, the Darmon points and the (weak) Campana points on $(\mathcal{X},\sum_{i\in I}\mathcal{D}_{i})$ are just the integral points on $\mathcal{X}\setminus\bigcup_{i\in I}\mathcal{D}_{i}$ as in Example 3.23.
@@
As explained in [30, §8], Darmon points can be viewed as integral points on the root stack $(\mathcal{X},\sqrt[\mathbf{m}]{\mathcal{D}})$ corresponding to the Campana pair $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$.
@@
Geometric Campana points can be viewed as the subset of Campana points which remain Campana points after enlarging the base field. In particular, geometric Campana points are just the same as Campana points whenever the support of the divisor $D_{\mathbf{m}}$ is a strict normal crossings divisor (over the ground field $K$). For Darmon points analogous statements hold.
Remark 3.28.
In this article, we only consider geometric Campana points and geometric Darmon points over rings of integers $\mathcal{O}_{S}$ which do not ramify in the splitting field $E/K$. In [45], geometric Campana points and geometric Darmon points of more general rings of integers were considered. In their setting, Campana points and Darmon points can differ from their geometric counterparts due to ramification, see [45, Lemma 2.7].
Let us consider an example where geometric Campana points differ from Campana points.
@@
Besides Campana points, Darmon points, and (weak) Campana points, there are many more interesting sets of $\mathcal{M}$-points, see [30, §3.3].
@@ 4. The Picard group of a pair and its geometry
For pairs $(X,M)$ corresponding to Darmon points, there is a natural definition for the Picard group of $(X,M)$. For such a pair, we have a corresponding root stack $(X,\sqrt[\mathbf{m}]{D})$ which lies over $X$. In this case we would like the Picard group of $(X,M)$ to be the Picard group of $(X,\sqrt[\mathbf{m}]{D})$, and the canonical divisor class to be the canonical divisor class on the stack. Both the Picard group and the canonical divisor class can be computed using the natural morphism $(X,\sqrt[\mathbf{m}]{D})\rightarrow X$, which is an isomorphism over the open set $U=X\setminus(D_{1}\cup\dots\cup D_{n})$ and ramified over the divisors $D_{i}$ with multiplicity $m_{i}$.
More generally, we can regard $(X,M)$ as a geometric space akin to a scheme or stack lying above $X$. If $(X,M)=(X,M_{\operatorname{mon}})$ then $(X,M)$ determines a functor $Y\mapsto(X,M)(Y)$ as we have seen in Remark 3.19. We view the natural inclusion map $(X,M)\rightarrow X$ as analogous to a “generically finite morphism” which is an isomorphism over $U$. Inspired by this view we introduce corresponding objects to the pair $(X,M)$, such as divisors and a Picard group.
@@ 4.1. Divisors on pairs
For a pair $(X,M)$ we define its set of generators to be the set $\Gamma_{M,\mathcal{C}}$ of all $(\mathbf{m},c)\in\mathfrak{M}_{\mathcal{C}}\cap\mathbb{N}^{n}_{\mathcal{C}}$ such that the only subset $J\subset\mathfrak{M}_{\mathcal{C}}\setminus\{(\mathbf{0},X)\}$ such that
@@
is $J=\{(\mathbf{m},c)\}$. If $(X,M)$ is a divisorial pair defined by a set of multiplicities $\mathfrak{M}\subset\overline{\mathbb{N}}^{n}$, then we set
@@
    every divisor $D_{i}$ is connected, nonempty and smooth over $K$ while $\sum_{i=1}^{n}D_{i}$ is a strict normal crossings divisor as defined in [48, Tag 0BI9],
The use of the term “smooth” is motivated by the fact that strict normal crossings pairs as considered in logarithmic geometry are log smooth, see e.g. [33, Chapter IV, Example 3.1.14]. Furthermore, [30, Proposition 8.3] shows that a root stack $(X,\sqrt[\mathbf{m}]{D})$ is smooth if its corresponding pair $(X,M)$ is smooth.
@@
where $c^{\prime}\subset c$. Then the pair $(X,M^{\prime})$ over $K$ defined using this data is equivalent to $(X,M)$. Furthermore, by the assumptions on $(X,M)$ it follows that $(X,M^{\prime})$ is smooth. ∎
@@
Let $D$ be a divisor on $X$ and let $(\mathbf{m},c)\in\mathbb{N}^{n}_{\mathcal{C}}\setminus\{(\mathbf{0},X)\}$. We define
$\mu((\mathbf{m},c),E-E^{\prime}):=\gcd(\mathbf{m})\cdot v_{(\mathbf{m}/\gcd(\mathbf{m}),c)}(f),$
where $f\in K(X)^{\times}$ is a local equation for $D$ on some open set intersecting $c$.
@@
The pullback of divisors to $(X,M)$ is intimately related to the theory of weighted (stacky) blow-ups as described in [40]. For $(\mathbf{m},c)\in\mathbb{N}_{\mathcal{C}}^{n}$, consider an open subset $X^{\prime}\subset X$ such that the divisors $D_{1},\dots,D_{n}$ are principal on $X^{\prime}$ and $\tilde{D}_{\mathbf{m}}=\tilde{D}_{\mathbf{m},c}$. Let $I_{\bullet}$ be the graded $\mathcal{O}_{X^{\prime}}$-subalgebra of $\mathcal{O}_{X^{\prime}}[t]$ generated by $f_{1}t^{m_{1}},\dots,f_{n}t^{m_{n}}$, where $f_{i}=0$ is a local equation for $D_{i}$. If $\pi\colon\mathrm{Bl}_{I_{\bullet}}X^{\prime}\rightarrow X^{\prime}$ is the corresponding weighted blow-up as defined in [40, Definition 3.2.1] and $D\in\operatorname{Div}(X^{\prime})$, then the coefficient of the exceptional divisor in $\pi^{*}D$ is equal to $\mu((\mathbf{m},c),D)$.
@@ 4.2. The Picard group of a pair
If $(X,M)$ is a smooth split pair corresponding to Darmon points with associated root stack $(X,\sqrt[\mathbf{m}]{D})$ as in [30, §8], then [9, Corollary 3.12] implies that the map $\operatorname{Pic}(X,M)\rightarrow\operatorname{Pic}(X,\sqrt[\mathbf{m}]{D})$, given by sending $\tilde{D}_{m_{i}\mathbf{e}}$ to the Cartier divisor $\frac{1}{m_{i}}D_{i}$ for all $i\in\{1,\dots,n\}$ with $m_{i}<\infty$, is an isomorphism. Furthermore, this isomorphism is compatible with the pullback homomorphisms of $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,M)$ and $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,\sqrt[\mathbf{m}]{D})$.
@@
Since $(X,M)$ is proper, for each $i=1,\dots,n$ there exists an element $(d_{i}\mathbf{e}_{i},D_{i})\in\Gamma_{M,\mathcal{C}}$ for some integer $d_{i}$. In particular, the coefficient of $\tilde{D}_{d_{i}\mathbf{e}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D_{i}$ is $d_{i}$, while for all $j\in\{1,\dots,n\}$ different from $i$, the coefficient of $\tilde{D}_{d_{i}\mathbf{e}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D_{j}$ is zero, as $D_{i}$ is not contained in $D_{j}$. Similarly, for all $D\in\operatorname{Div}(U)$, the the coefficient of $\tilde{D}_{d_{i}\mathbf{e}_{i},D_{i}}$ in $\operatorname{pr}_{M}^{*}D$ is also zero. Since, furthermore, the restriction of $\operatorname{pr}_{M}^{*}$ to $\operatorname{Div}(U)$ is injective, $\operatorname{pr}^{*}_{M}$ is indeed injective. This directly implies that the homomorphism on the Picard groups is also injective.
@@
Suppose that $D\in\operatorname{Div}(X,M)$ is an element whose image in $\operatorname{Pic}(X_{\overline{K}},M)$ is trivial. This implies that $D$ is the pullback of $\operatorname{div}f$ for $f\in\overline{K}(X)^{\times}$. The rational function $f$ is defined over some Galois extension $E/K$, and $\tilde{f}=\prod_{\sigma\in\operatorname{Gal}(E/K)}\sigma(f)\in K(X)^{\times}$. Since $D$ is invariant under the Galois action it follows that the pullback of $\operatorname{div}(\sigma(f))$ to $(X_{E},M)$ is $D$ itself. Therefore, it follows that $\deg(E/K)D\in\operatorname{Div}(X,M)$ is a principal divisor, so $D]\in\operatorname{Pic}(X,M)$ is torsion, as desired. ∎
@@ 4.3. The canonical divisor and the effective cone
Thus the canonical class is defined in such a way that it satisfies an analogue of the Hurwitz formula for morphisms of curves (see e.g. [23, Chapter IV, Proposition 2.3]). If $(X,M)$ is a smooth pair corresponding to Darmon points on a Campana pair $(X,D_{\mathbf{m}})$, then the canonical divisor of $(X,M)$ agrees with the canonical divisor of the root stack $(X,\sqrt[\mathbf{m}]{D})$, see for example [51, Proposition 5.5.6] for the case when $X$ is a curve.
@@
The effective cone of a proper pair is strictly convex, as the next proposition shows.
@@
In fact, for computing the Fujita invariant and the $b$-invariant, we only need to consider small generators $(\mathbf{m},c)\in\Gamma_{M,\mathcal{C}}$, which often considerably simplifies computations. In particular, we only need to consider the minimal elements for the partial ordering on $\Gamma_{M,\mathcal{C}}$ given by $(\mathbf{m},c)\leq(\mathbf{m}^{\prime},c^{\prime})$ if $\mathbf{m}\leq\mathbf{m}^{\prime}$ and $c\supset c^{\prime}$. The following lemma shows that we need to consider even less elements in $\Gamma_{M,\mathcal{C}}$. For a stratum $c$ we set
@@
Note that here it suffices to consider the smallest strata for the polyhedra, since if $c\subset c^{\prime}$, then $P_{c^{\prime}}\subset P_{c}$. Furthermore, for a divisorial pair $(X,M)$, the pairs $(X,M^{\prime})$ and $(X,M^{\prime\prime})$ are both divisorial, with sets of multiplicities $\mathfrak{M}^{\prime}=\Gamma_{M}\cap\partial P$ and $\mathfrak{M}^{\prime\prime}=\Gamma_{M}\cap V(P)$, where $P$ is the polyhedron given as the convex hull of the set
@@
This is essentially the same argument as in Lemma 4.39 by using Lemma 4.21. ∎
@@ 5. Rationally connected pairs
In modern times, Manin’s Conjecture is often formulated for rationally connected varieties, see for example [27, Conjecture 1.2]. The notion of rationally connected varieties (see e.g. [26, Definition-Remark 2.2]) has a natural extension to smooth proper pairs.
@@
A smooth proper pair $(X,M)$ over a field $K$ is rationally connected if there exists a nonempty open subvariety $V\subset X$ such that for each algebraically closed field $L/K$ and every two points $p_{1},p_{2}\in V(L)$, there exists a rational curve $C\subset X_{L}$ containing both points such that $C$ is the image of a morphism $f\in(X,M_{\operatorname{mon}})(\mathbb{P}^{1}_{L})$, where $(X,M_{\operatorname{mon}})$ is the pair from Notation 3.18.
@@
By spreading out [39, Theorem 3.2.1], we can find a set of places $S^{\prime}$ containing $S$ such that $f\colon\mathbb{P}^{1}_{K}\rightarrow X$ lifts to a morphism $\tilde{f}\colon\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}\rightarrow\mathcal{X}$. The pullbacks $\tilde{f}^{*}\mathcal{D}_{1},\dots,\tilde{f}^{*}\mathcal{D}_{n}$ are effective divisors on $\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}$. By further enlarging $S^{\prime}$ if necessary, we can ensure that these divisors have no components supported above a prime $\mathfrak{p}$ in $\mathcal{O}_{S^{\prime}}$, which ensures that $\tilde{f}\in(\mathcal{X},\mathcal{M}_{\operatorname{mon}})(\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}})$. Let $P\in\mathbb{P}_{K}^{1}(K)$. By the valuative criterion of properness, $P$ corresponds to an unique integral point $\tilde{P}\colon\operatorname{Spec}\mathcal{O}_{S^{\prime}}\rightarrow\mathbb{P}^{1}_{\mathcal{O}_{S^{\prime}}}$. Now it follows from the functoriality discussed in Remark 3.19 that $\tilde{f}\circ\tilde{P}\in(\mathcal{X},\mathcal{M}_{\operatorname{mon}})(\mathcal{O}_{S^{\prime}})$. ∎
@@
Let $(X,M)$ be a smooth proper pair over a field $K$ such that $X$ is a toric variety and $D_{1},\dots,D_{n}$ are torus-invariant prime divisors. Then $(X,M)$ is rationally connected.
@@
Without loss of generality, we can assume that $\mathfrak{M}$ is a monoid and that $X$ is a split toric variety (so that $(X,M)$ is a split pair). Let $L$ be an algebraically closed field containing $K$ and let $L(t)$ be the field of rational functions over $L$ in the variable $t$. Write $(X_{L(t)},M_{L(t)})$ for the pair over $L(t)$ obtained by base changing $X$ and $D_{1},\dots,D_{n}$ to $L(t)$. This pair has the obvious integral model $(X_{\mathbb{P}^{1}_{L}},M_{\mathbb{P}^{1}_{L}})$ over $\mathbb{P}^{1}_{L}$ given by base changing $(X,M)$ to $\mathbb{P}^{1}_{L}$. Now [30, Theorem 1.3] implies that the pair $(X_{L(t)},M_{L(t)})$ satisfies $M_{L(t)}$-approximation off the place $(1:0)\in\mathbb{P}_{L}^{1}$. Consequently, the embedding